OpenStudy (anonymous):

5x - y = 8 Part A- Explain how you will show all of the solutions that satisfy this equation. Part B- determine 3 different solutions for this equation. Part C- Write an equation that can be paired with the given equation in order to form a system of equations that is inconsistent. @undeadknight26

OpenStudy (anonymous):

@phi @SolomonZelman

zepdrix (zepdrix):

Hey Brie :) Mmmm I'm not exactly sure what they're asking for in part A, maybe they want us to write it in terms of y= stuff. Let's start with part B though. Do you understand how to find solutions? Like if you let x=0, 5x - y = 8 becomes, 0 - y = 8 What y value does that give us?

OpenStudy (wwwhhhaaattt?):

@StudyGurl14

OpenStudy (anonymous):

i prefer Alex lol and no im so freakign tired and done with this test i cant even see straight can you dumb it up for me please?

OpenStudy (anonymous):

oh 8 right?

zepdrix (zepdrix):

This is given to us in the `standard form of a line`,\[\Large\rm 5x - y = 8\] So we have a line :) We can pick any point for one of our letters, and figure out what the other point is. So by choosing x=0,\[\Large\rm 5\cdot 0-y=8\]\[\Large\rm -y=8\]Multiply both sides by -1,\[\Large\rm y=-8\] This is telling us that one solution to the given equation is \(\Large\rm (0,-8)\)

zepdrix (zepdrix):

8 is close though ^^ hehe

zepdrix (zepdrix):

If we choose to let y=0, this will lead us to another solution. (zeros are easy to work with, so they're nice numbers to choose.)

OpenStudy (anonymous):

okieeeeee you have my full attention

zepdrix (zepdrix):

\[\Large\rm 5x-y=8\]\[\Large\rm 5x-0=8\]\[\Large\rm 5x=8\]Divide by 5, giving us,\[\Large\rm x=\frac{8}{5}\]So that gives us our next solution, \(\Large\rm \left(\dfrac{8}{5},0\right)\).

zepdrix (zepdrix):

Realize that this is a line, there will be an infinite number of solutions. We're just choosing the ones that are easy.

OpenStudy (anonymous):

ok and then there is one more right?

zepdrix (zepdrix):

There are infinitely many more. But yes we need just one more :) How about if x=2? Can you figure out the y that goes along with that?

OpenStudy (anonymous):

oh lordy let me try

OpenStudy (anonymous):

10-y=8? then y would equal 6?

zepdrix (zepdrix):

10 - 6 = 8 ? 0_o Oh boy, you really are tired, huh?

OpenStudy (anonymous):

ya

zepdrix (zepdrix):

How far apart are 10 and 8? -_- If you subtract 6 from 10, you end up at 4, not at 8 like we need.

OpenStudy (anonymous):

sorry y would be 12

zepdrix (zepdrix):

whut? :U brie brie brie.....

zepdrix (zepdrix):

You're at 10.... you need to subtract some amount to get to 8. So you would subtract.... 2, yes? 10 minus 2 is 8 Yes? 0_o

OpenStudy (anonymous):

omg I told you I was tired I cant think sorry

zepdrix (zepdrix):

So when we chose x=2, we ended up with a y value of 2 as well. Do you understand how to write that as a coordinate pair?

zepdrix (zepdrix):

When x=0, we found our y to be -8. And we wrote it as (0,-8). How bout this third solution? How we gonna write it? :U

OpenStudy (anonymous):

I really don tknow im so sorry

zepdrix (zepdrix):

Mmmm k, moving on :U For Part C: In order for a system to be `incosistent`, it means that the system has `no solutions`. This happens when lines are `parallel` but have different `y-intercepts`. \(\Large\rm 5x-y=8\) We'll multiply our x and y by some amount, then throw a different number on the right side. If we multiplied by 2,\[\Large\rm 10x-2y=16\]And throw a different number on the right side,\[\Large\rm 10x-2y=10\]This gives us, with the first equation we were given, a system this is inconsistent. \[\Large\rm inconsistent\quad \cases{5x-y&=8\\ \rm 10x-2y&=10}\]

zepdrix (zepdrix):

Ok your brain must be fried by now. Go eat some pringles, then go to bed! :)

OpenStudy (anonymous):

pringles?

zepdrix (zepdrix):

Yah, it's a brand of... chips.. or whatever XD

OpenStudy (anonymous):

so wait what was the last solution ? @zepdrix

OpenStudy (anonymous):

WOW you reply right after I submit my exam -_-

zepdrix (zepdrix):

Our last solution? :U Comeoooooon, don't be a slacker. You gotta contribute something here :( Our first solution was given by x=0, y=-8, so we wrote our coordinate pair (x,y) as (0,-8). Our second solution was x=8/5, y=0, so we wrote our coordinate pair (x,y) as (8/5, 0). Our third solution was x=2, y=2, so we would write our coordinate pair (x,y) as ?

zepdrix (zepdrix):

You have to contribute something brie :( I can't just do the entire assignment for you..